# Visualizing Causality in Mixed Reality for Manual Task Learning: A Study

### v3

@@ VIII-A Future Work
VIII-A1 Dynamic Causality
As described in  subsection VII-D, showing merely one future causally connected to the present results in neglected details is not optimal for learning. Therefore, we propose enhancing the representation of causality and human intention in manual task learning by dynamically modeling causality. In other words, the learner sees how the causally connected futures are affected by their actions in the present. Furthermore, by showing the future dynamically, the learners can visualize and understand a broader range of causality, providing a complete understanding of the task.
@@ VIII-A1 Dynamic Causality
VIII-A2 Visualization of Causality
One of the less satisfying results from our study is the compromise in learning time learners have to make when learning additional information about causality and intention of the task, as mentioned in  subsection VII-C. There is always a trade-off between the content’s volume and the time spent learning. For example, ”I could understand that my action would reflect the task, but it took me a while to realize how (G4, P11.)” Therefore, we suggest that natural, user-friendly, and dedicated visualization techniques can be applied to causality and intention in MR. Take the scenario of making an omelet for example. Should the learners set the heat too high when preheating the pan, a 3D animation of an omelet being overcooked can be overlaid on the pan in MR. Such dedicated visualizations can intuitively acknowledge the user’s causality within the tasks without significantly increasing the time and cognitive cost. Yet, such visualizations’ design, formulation, and generalization remain challenging. One promising direction is to utilize human attention during an action. State-of-the-art artificial intelligence algorithms can predict human attention based on their actions  [105, 106, 107]. By predicting the engagement based on learners’ interaction with the objects and better positioning the visualization, it is possible that the learner can naturally perceive knowledge without a high cognitive cost in transitioning.
VIII-A3 Adaptive Causality
Another workaround to reduce the time and cognitive cost of learning is applying adaptive Levels of Detail in the causality and intention of the learners. As suggested in  [89, 90, 91], learners can infer the demonstrator’s causality and intention while observing. Our findings, to some degree, suggest that even without explicitly presenting the causality and intention, the learners can infer them (Understanding - Q1,2,3). We identify this ability as active inference in  subsection VII-B. Based on our work, we anticipate future research on two major research questions:
How much of the causality can learners actively infer? In our study, we conclude that learners cannot actively infer all interaction-level causality. Yet, how much of the causality can learners infer? And how do individual intelligence and cognition contribute to the inference?
What is the difference between the preserved and the inferred causality? A promising research prospect is to detect the learners’ inference of causality and to quantify and measure the distance between the preserved causality from authors and the inferred causality from learners. We envisage that future research will shed light on quantifying the understanding of causality in MR manual task learning.
With knowledge of the inferred causality of the learners, adaptive learning methods can be applied. Should more causal information be provided to learners who infer less and vice versa, and intended causality supplied to learners who infer differently, we expect more efficiency and less cognitive load in MR manual task learning methods.
@@ Acknowledgement
This work is partially supported by the NSF under the Future of Work at the Human-Technology Frontier (FW-HTF) 1839971. We also acknowledge the Feddersen Distinguished Professorship Funds and a gift from Thomas J. Malott. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the funding agency.

### v4

@@ VIII-A Future Work
VIII-A1 Visualization of Causality
One of the less satisfying results from our study is the compromise in learning time learners have to make when learning additional information about causality and intention of the task, as mentioned in  subsection VII-B. There is always a trade-off between the content’s volume and the time spent learning. For example, ”I could understand that my action would reflect the task, but it took me a while to realize how (G4, P11.)” Therefore, we suggest that natural, user-friendly, and dedicated visualization techniques can be applied to causality and intention in MR. Take the scenario of making an omelet for example. Should the learners set the heat too high when preheating the pan, a 3D animation of an omelet being overcooked can be overlaid on the pan in MR. Such dedicated visualizations can intuitively acknowledge the user’s causality within the tasks without significantly increasing the time and cognitive cost. Yet, such visualizations’ design, formulation, and generalization remain challenging. One promising direction is to utilize human attention during an action. State-of-the-art artificial intelligence algorithms can predict human attention based on their actions  [106, 107, 108]. By predicting the engagement based on learners’ interaction with the objects and better positioning the visualization, it is possible that the learner can naturally perceive knowledge without a high cognitive cost in transitioning.
VIII-A2 Dynamic Causality
Showing merely one future causally connected to the present results in neglected details is not optimal for learning. Therefore, we propose enhancing the representation of causality and human intention in manual task learning by dynamically modeling causality. In other words, the learner sees how the causally connected futures are affected by their actions in the present. Furthermore, by showing the future dynamically, the learners can visualize and understand a broader range of causality, providing a complete understanding of the task.
@@ VIII-A2 Dynamic Causality
VIII-A3 Comparing methods for task decomposition
In our study, the tasks are decomposed based on three levels of causal hierarchy derived from cognitive psychology [27]. We employed an object-level segmentation of tasks, where causality is determined based on the relationships between tasks that share common objects  [25]. While we focus on identifying and categorizing the levels of causality in manual tasks and design experiments specifically for causality, we acknowledge that other forms of task decomposition methods can be used to segment the task and influence the learning results. Vitally, prior work [111] has shown that the information shown in each visualization has a great impact on the performance of the users following the visualization. We envision future works to incorporate our study findings into state-of-the-art task decomposition methods to optimize task learning with interactive systems [112].
VIII-A4 Causality addressing Error Cascades in Task Execution
Initial errors can lead to a chain of mistakes which will impact subsequent task performance and reduce the overall performance. Visualizing causal links within the task can help users understand how a current error influences future steps and the final outcome of the task. This type of error visualization highlights why each step is important and how it contributes to the overarching goal of task completion. Future directions include implementing real-time error detection methods, employing adaptive visualizations to illustrate the cascading effects of errors on subsequent actions, and providing corrective feedback. These can improve task learning by mitigating the negative impact of errors, enhance understanding, and improve learning gains.
@@ VIII-B Limitation
VIII-B1 Pre-test for knowledge gathering
Our study considered only post-test knowledge testing with a two-fold reasoning. First, participants were new to the task and unfamiliar with our MR system of causality, and hence we assumed a uniform baseline of knowledge. Second, considering the manual task procedure and operations involved in our study are monotonous, we forwent a pre-test to avoid potential learning effects that could mitigate the findings in our post-test. Therefore, we acknowledge that an ideal pre-test could have been conducted to assess the participants’ prior knowledge of the manual tasks in our study, and therefore could have constituted a more comprehensive analysis of the learning gain.
VIII-B2 Test-bed Limitations
@@
VIII-B3 User Study Sample Size
Following the convention of 12 participants per group for usability and quantitative evaluation in HCI [114] , we included 12 participants per condition in our study. We analyzed the effect sizes post hoc. We found that our sample size of 12 produced varying levels of effect sizes from small effect to large effect. We specifically found large effect sizes for metrics such as learning time, number of repeats, and completion time in the testing phase. We found small effects for the mistake calculation. For NASA TLX and subjective feedback, the effect again varies. As we have results for mixed effect sizes, sample size may have been small to detect smaller or medium effect sizes reliably. While the current sample size was sufficient for detecting the observed differences in large effect size, increasing the number of participants could further improve the precision of effect size estimates and allow for the detection of small or medium effect. This adjustment will provide greater statistical power, reduce variability, and confirm the reliability of observed effects.
@@ Acknowledgement
This work is partially supported by the NSF under the Future of Work at the Human-Technology Frontier (FW-HTF) 1839971 and NSF Partnerships for Innovation Technology Transfer (PFI-TT) 2329804. We also acknowledge the Feddersen Distinguished Professorship Funds and a gift from Thomas J. Malott. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the funding agency.